OpenStudy (richyw):

Inverse Fourier transform of \(F(w)=e^{-|w|a}\) where a > 0

OpenStudy (richyw):

\[f(x)=\int^\infty_{-\infty}e^{-iwx}e^{-|w|a}dw\]\[f(x)=\int^0_{-\infty}e^{w(a-ix)}dw+\int^\infty_0e^{w(-a-ix)}dw\]The integral on the right is messing me up. I make the substitution \[u=w(a-ix)\]\[du=(a-ix)dw\]

OpenStudy (richyw):

oops, that's the substitution for the first integral. I substitute \[v = w(-a-ix)\]\[dv=(-a-ix)dw\]

OpenStudy (richyw):

\[\frac{1}{-a-ix}\int_0^{?} e^vdv\]

OpenStudy (richyw):

is my top limit of integration now \(-\infty\) ? because \(v=-w(a+ix)\) If not I would get an infinite answer...

OpenStudy (richyw):

ok I think I just missed that detail about the integration limits in the integral. I end up with \[f(x)=\frac{2a}{a^2+x^2}\]

OpenStudy (accessdenied):

At the beginning when you take the inverse Fourier transform by integral definition, you have to use the positive exponent right? \( \large e^{i \omega x} \) instead of \( \large e^{-i \omega x} \)

OpenStudy (accessdenied):

But it seems like I get the same answer with that adjustment.

OpenStudy (richyw):

doesn't it depend on how you define the ft and ivt? like one of them has to have the 2π factor and the other one doesn't.

OpenStudy (accessdenied):

Hm... I knew there was some uncertainty around the 2 pi factor for definition, but I actually did not know there was room to define the e^(iwx) vs. e^(-iwx) differently as well. :) http://reference.wolfram.com/mathematica/tutorial/IntegralTransformsAndRelatedOperations.html

OpenStudy (accessdenied):

So, I figure as long as that is consistent with what you use normally, that will be fine then. It seems both led to the answer.

OpenStudy (richyw):

yup, that page you have even shows the different conventions. in physics you usually see it different. Apparently this is the "pure mathematics" convention. funny because the book is called "applied PDE's" haha. Also thanks for sending that link, because now I know how to change the convention from the default in mathematica!

OpenStudy (accessdenied):

yup, I found it looking up the situation more closely. Linked through a stackexchange post about it, top response. http://math.stackexchange.com/questions/574581/fourier-transform-convention-frac1-sqrt2-pi-int-infty-infty-fx?rq=1